6/15d+9/30d-1/6d=d+1.5-2/12d
1
Expert's answer
2018-04-07T04:44:11-0400
d(6/15 + 3/10 - 1/6 - 1 + 1/6) = 3/2
d(6/15 + 3/10 - 1) = 3/2
d(12 + 9 - 30)/30 = 3/2
d = - 3*30/(2*9) = - 5